# Enhancing Diffusion Models Towards Unified Generative and Discriminative Learning

### v2

@@ 5.2 Performance and scalability on ImageNet
FID-50k; denoted as w/ CFG – w/o CFG for conditional models.
@@
Comparison with advanced diffusion systems. Our baseline on Lightning-DiT (lightning) is already highly-optimized and, with REPA (repa), represents the state-of-the-art with extremely fast convergence. Against this strong baseline, it is compelling that our minimal tweak yields further gains. Note that while REPA achieves high accuracy (69.4%) via direct DINOv2 alignment, this gain does not fully translate into a superior FID (10.0, worse than Lightning-DiT and ours), a disconnect similar to Sec. 5.1. This suggests that externally boosting feature quality alone is insufficient, if the internal decoder remains architecturally decoupled from discriminative gains. We identify this missing link and explore an alternative ab initio way, aiming to structurally bridge it.
@@ Appendix B Implementation details
ViT-based architecture UViT-S DiT-B DiT-L
patch size 2 1 1
hidden size 512 768 1024
ViT layers 13 12 24
attention heads 8 12 16
(param count) 44.3M 129.8M 457.1M
@@
  • Semantic segmentation (evaluation-time): Following iBOT (ibot), to leverage multi-level features for higher performance, we extract patch tokens from two fixed intermediate layers (5th and 7th, for both baselines and ours) and concatenate them. The optimal timesteps obtained by searching are 0.25 (ADE20K) and 0.45 (VOC), slightly noisier than the one in classification.
@@ Appendix F Limitations and future research directions
Although we conduct extensive experiments to prove our claim that it is possible to enhance both generative and discriminative performance simultaneously through our approaches, our evaluations are limited to datasets up to ImageNet-1k ($256^{2}$) and models up to DiT-L. We did not investigate larger datasets with higher resolutions and larger models, or even on text-to-image generation, due to high computational costs inherently for generative models.

### v3

@@ 5.2 Performance and scalability on ImageNet
FID-50k; denoted as (w/ CFG –) w/o CFG for conditional models.
@@
Comparison with advanced diffusion systems. Our baseline on Lightning-DiT (lightning) is already highly-optimized and, with REPA (repa), represents the state-of-the-art with extremely fast convergence. Against this strong baseline, it is compelling that our minimal tweak yields further gains. Note that while REPA achieves high accuracy via direct DINOv2 alignment, this gain does not fully translate into a superior FID, a disconnect similar to Sec. 5.1. This suggests that externally boosting feature quality alone is insufficient, if the internal decoder remains architecturally decoupled from discriminative gains. We identify this missing link and explore an alternative ab initio way, aiming to structurally bridge it.
@@ Appendix B Implementation details
ViT-based architecture UViT-S DiT-B DiT-L DiT-XL
patch size 2 1 1 1
hidden size 512 768 1024 1152
ViT layers 13 12 24 28
attention heads 8 12 16 16
(param count) 44.3M 130M 457M 675M
@@
  • Semantic segmentation (evaluation-time): Following iBOT (ibot), to leverage multi-level features for higher performance, we extract patch tokens from multiple intermediate layers ({5, 7}-th for DiT-B and {9, 11, 13, 15}-th for DiT-L, fixed for both baseline and our models) and concatenate them. The optimal timesteps obtained by searching are 0.25 (ADE20K) and 0.45 (VOC2012), slightly noisier than the one used in classification.
@@ Appendix F Limitations and future research directions
Although we conduct extensive experiments to prove our claim that it is possible to enhance both generative and discriminative performance simultaneously through our approaches, our evaluations are limited to datasets up to ImageNet-1k ($256^{2}$). We did not investigate larger datasets with higher resolutions and larger models, or even on text-to-image generation, due to high computational costs inherently for generative models.